A car travelling south at 36 m/s sounds a 456 Hz horn.

(a) What frequency is heard by the driver in an approaching truck traveling north at 25 m/s?

(b) What frequency is heard by the driver if he is heading south away from the car at 25 m/s?


1
Expert's answer
2021-12-02T13:58:49-0500

(a) From the Doppler Shift formula


"f=f_s\cdot\frac{v+v_0}{v-v_s}=456\cdot\frac{340+25}{340-36}=547.5\ (Hz)"


"v" - is the speed of the sound;

"v_0" - is the speed of the observer;

"v_s" - is the speed of the source;

"f_s" - is the source frequency.


(b) From the Doppler Shift formula


"f=f_s\cdot\frac{v-v_0}{v-v_s}=456\cdot\frac{340-25}{340-36}=572.5\ (Hz)"